can someone please explain to me why
4 sq root 16^3 is equal to 2^3 which is equal to 8
thank you

Answer by Alan3354(69443)

can someone please explain to me why
4 sq root 16^3 is equal to 2^3 which is equal to 8
-------------------
Who said it is?
4*sqrt(16^3) = 4*4^3 = 256
------------
That's 2^8, = 2^(2^3)